OpenStudy (anonymous):

What is the value of x in the equation 8x – 2y = 48, when y = 4?

OpenStudy (anonymous):

8x - 2y = 48 y = 4 Plug in '4' for 'y': 8x - 2(4) = 48 Now multiply -2 * 4

OpenStudy (anonymous):

\[8x-2(4)=48\]

OpenStudy (anonymous):

what u get when u multiply -2 with 4? @Rafyy